OpenStudy (donnie1999):

find the quotient -30s^3t-20st/10st

OpenStudy (anonymous):

Deez nutz

OpenStudy (anonymous):

LOL no comment

OpenStudy (anonymous):

i think but just to be sure @mathmath333 (−30s3)(t)−(20st10s)(t) =−30s3t+−2s2t2 Answer: =−30s3t−2s2t2

OpenStudy (anonymous):

Which value of c from the set {4, 5, 6, 7, 8} makes this equation true? 2c + 9 = 25 4 5 6 7 8

OpenStudy (anonymous):

8

OpenStudy (anonymous):

2c+9=25 Step 1: Subtract 9 from both sides. 2c+9−9=25−9 2c=16 Step 2: Divide both sides by 2. 2c2=162 c=8 Answer: c=8

OpenStudy (anonymous):

Thanks.